OpenStudy (anonymous):

what is the simplified form of (12y^4/8x^3) times (6x^8/20 y^6)

OpenStudy (anonymous):

\[\huge \frac {12y^{2}}{8x^{3}}*\frac{6x^{8}}{y^{6}}\] ????

OpenStudy (anonymous):

12y to the 4th power

OpenStudy (anonymous):

and 20 y to the 6th

OpenStudy (anonymous):

\[\huge \frac{12y^{4}}{8x^{3}}*\frac {6x^8}{20y^{6}}\] \[\huge \frac{3y^{4}}{4x^{3}}*\frac {3x^8}{5y^{6}}\] Now cancel the powers.

OpenStudy (anonymous):

\[\huge \frac{3}{4}*\frac {3x^5}{5y^{2}}\] \[\huge \frac {9x^5}{20y^{2}}\]